# Susan Athey    Emil Palikot

### v4

@@ 6.3 Counterfactual Mentors’ Selection Policies
Applicants selected by mentors and randomly assigned to the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors identify promising applicants who are likely to secure a tech job regardless of their participation in Mentoring. Table 8 compares characteristics of selected and non-selected applicants (panel A), and selected by different groups of mentors (panels B and C). Baseline is the share of control group applicants that found a job.
@@
Panel A of Table 8 reveals systematic differences between applicants selected by mentors and those not selected. Selected applicants exhibit substantially stronger baseline employment prospects (0.456 vs 0.286), despite having similar years of professional experience (7.1 vs 7.5 years). Selected applicants are more likely to be over 30 years old (0.487 vs 0.147) and to hold STEM degrees (0.510 vs 0.182), while being considerably less likely to come from small towns (0.245 vs 0.476). They are also less likely to have social science backgrounds (0.134 vs 0.207).
@@
Panel B compares mentors with versus without managerial experience. Mentors with managerial experience select applicants with weaker baseline employment prospects (0.378 vs 0.549) but who are substantially older (0.588 vs 0.378) and have more professional experience (7.6 vs 6.5 years). Panel C shows that mentors with above-median total career experience exhibit similar but more pronounced patterns. They select applicants with weaker baselines (0.390 vs 0.512) who are older (0.590 vs 0.419), have substantially more professional experience (8.2 vs 6.3 years), and are more likely to hold a STEM degree (0.564 vs 0.480). The consistency across Panels B and C suggests a common selection philosophy among more experienced mentors: they look beyond current employability to identify "high-potential, high-barrier". Less experienced mentors appear to select more conventionally, favoring applicants already demonstrating stronger baseline employability.
We use our off-policy tools to examine alternative selection strategies that mentors could adopt and their potential implications for the mentoring program’s effectiveness. Specifically, we consider four scenarios: (i) Highest CATE—targeting applicants predicted to have the highest treatment effects, (ii) Most Promising—targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising—selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated—selecting participants with the highest predicted outcomes if they participate in Mentoring. Additionally, we include two benchmarks: mentors selecting at random (Random Choice) and the observed allocation (Observed). In each of the four policies, we assume that 15% of applicants participate in Mentoring, which is an approximate share of participants to applicants observed by DareIT in recent years.
Table 9 examines the extent to which different allocation policies select the same individuals Each cell shows the percentage of individuals in the row policy who also appear in the column policy. First, we consider the relationship between mentor-based selection (Current) and the Highest CATE algorithmic policy. Among current mentor-selected participants, 70.4% would also be assigned to Mentoring under the Highest CATE algorithm, demonstrating substantial concordance between expert judgment and data-driven targeting (note that we do not use the information on whether a candidate was selected or not in the policy estimation).
Current mentor selections align most closely with Highest CATE (70.4%), followed by Most Promising (29.6%), with little overlap with Least Promising (18.3%) or Highest Chances (19.7%). This pattern suggests mentors implicitly optimize for treatment effects rather than baseline characteristics or final outcomes, though with some cream-skimming tendency. The 29.6% overlap with Most Promising indicates approximately one-third of mentor selections favor strong baseline candidates, while the predominant 70% overlap with Highest CATE confirms that mentors primarily target based on perceived program responsiveness. The low overlap with Least Promising (18.3%) suggests mentors, whether by design or implicit bias, rarely select individuals with the weakest baseline labor market attachment, even when such individuals might benefit from treatment.
Table 9: Overlap Between Allocation Policies
Policy H. CATE Most Prom. Least Prom. H. Chances Current
Highest CATE 100.0
@@
Highest Chances 27.9 26.7 31.5 100.0
Current 70.4 29.6 18.3 19.7 100.0
Notes: Each cell shows the percentage of individuals in the row policy that also appear in the column policy. Only lower triangle shown to avoid redundancy; upper triangle omitted. Sample sizes in parentheses after policy names. Diagonal entries are 100% by construction (bold). H. CATE = Highest CATE; H. Chances = Highest Chances; Most/Least Prom. = Most/Least Promising.
Table 10 compares treatment effects across six allocation policies for the Mentoring program. The results demonstrate that targeting based on predicted treatment effects substantially outperforms alternative strategies. The Highest CATE policy achieves treatment effects of 3.59-3.59 months (SE = 0.033), the largest magnitude across all policies. Random allocation serves as the baseline benchmark, generating treatment effects of 2.20-2.20 months (SE = 0.040). Current mentor-based selection performs remarkably well, achieving 3.25-3.25 months (SE = 0.100), a 48% improvement over random allocation. This validates that mentors successfully identify high-benefit participants through their evaluative selection process.
Table 10: Allocation Policy Comparison: Mentoring Program
Policy Baseline Time Treatment Effect
@@
Data-driven policies:
Highest CATE 11.08 (0.086) -3.59 (0.033)
@@
Highest Chances 11.77 (0.152) -2.14 (0.079)
Notes: Comparison of allocation policies for the Mentoring program. Baseline Time is the predicted months to employment without treatment (lower is better); Treatment Effect is the predicted change in months from treatment (more negative = faster employment). Standard errors in parentheses. Current reflects mentor-selected participants. Random is random assignment benchmark. Highest CATE maximizes treatment effects; Most/Least Promising target best/worst baseline outcomes; Highest Chances minimizes post-treatment time to employment.
Notably, the Highest CATE algorithmic policy achieves an additional 10.4% improvement over mentor selection, increasing treatment effects from 3.25-3.25 to 3.59-3.59 months. We obtain the 0.34-month gain using only six readily-observable characteristics (education, location, age, experience) while mentors access rich soft information from interviews and holistic assessment. The overlap analysis (Table 9) reveals that 70.4% of mentor-selected participants would also be assigned under Highest CATE, demonstrating substantial concordance between expert judgment and data-driven targeting. The finding that simple observable characteristics can meaningfully improve upon expert selection—even when experts demonstrate strong performance—underscores the value of systematic data-driven approaches as complements to human judgment.
Policies optimizing objectives other than treatment effects yield substantially weaker results. Most Promising, which targets applicants with the strongest baseline prospects (9.71 months to employment), achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation despite successfully cream-skimming the highest-potential candidates. This demonstrates that strong baseline prospects do not predict strong treatment responsiveness. Conversely, Least Promising targets individuals with weakest baselines (13.44 months) but generates the smallest treatment effects (1.94-1.94 months), revealing an unfortunate equity-efficiency tradeoff: those who need help most benefit least from the program. Highest Chances, which optimizes final employment outcomes, achieves 2.14-2.14 months, marginally worse than random assignment. These comparisons underscore that maximizing program impact requires explicitly targeting predicted treatment effects rather than baseline characteristics or final outcomes, as treatment response heterogeneity does not align with baseline heterogeneity.
@@ 7 Conclusion
Changing labor markets push workers to want to transition into high-growth sectors, and facilitating these transitions is an important goal for labor policy. However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity. In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector. This way we identify a class of interventions that can effectively narrow the gender gap in technology.
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by over 30% twelve months after the end of the program. We estimate that Mentoring increases technology employment from 46% to 61%, while Challenges increases it from 27% to 38%, with treatment effects persisting over the full observation period. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring. These differential patterns provide evidence on the underlying mechanisms: Mentoring expands professional networks (increasing LinkedIn connections by 52.7, a 23% increase), while Challenges provides credible skill signals. Notably, network expansion occurs only in Mentoring, and heterogeneity patterns align with each program’s mechanism: those with weaker networks benefit most from Mentoring, while those with weaker credentials benefit most from Challenges.
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate that optimal targeting achieves outcomes 86% larger than random assignment at current capacity levels (30% of applicants). Notably, the value of targeting persists even as capacity expands: at 50% capacity per program, targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that targeting generates value both through selecting who participates and through matching individuals to their highest-benefit program. The complementarity between individual characteristics and program type creates allocation gains beyond simple capacity expansion.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. Comparing selected-but-control applicants to non-selected applicants reveals that mentors successfully identify promising candidates: mentor-selected participants achieve outcomes 48% larger than random allocation. However, mentors appear to prioritize candidates with strong baseline prospects, those likely to secure technology employment even without the program, rather than those who would benefit most from participation. We evaluate an alternative algorithmic allocation policy that selects participants based on predicted treatment effects using only six readily-observable characteristics (education, location, age, experience). Despite mentors’ rich soft information from interviews and holistic assessment, algorithmic targeting achieves outcomes 63% larger than random allocation, representing an additional 11% improvement over mentor judgment. This finding is particularly striking given that algorithmic and mentor-based selections exhibit 70% concordance: the algorithm largely validates expert choices while identifying additional high-benefit candidates. The results demonstrate that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity, and that simple demographic targeting can complement human judgment to enhance program effectiveness.
This study comes with certain limitations worth noting. First, the limited participant size in both programs led to wide confidence intervals in several estimates, leaving room for uncertainty regarding the exact impacts of Mentoring and Challenges. Secondly, we could not pinpoint the exact mechanisms driving the outcomes for both programs. Although our findings suggest that Mentoring likely enhances participants’ professional networks and teaches how to utilize them, and Challenges supports participants in the development of a credible signal of their skills, a more in-depth exploration is required to understand the specific mechanisms through which the training provided by both programs impacts employability. Such an analysis could also help scale up Mentoring by identifying the crucial aspects of the program and directing Dare IT resources to them. Lastly, the participants from the first edition of Challenges may not be representative of those in later editions. Early adopters of such programs might experience greater benefits. To validate the program’s continued efficacy, subsequent evaluations are essential.
@@ Appendix B Covariate balance
In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only variable that has a statistically significant differnece is social science.
@@
Appendix C Kaplan Meier Survival Curves
@@ Appendix E Outcomes measured in surveys
We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes. There were three surveys: two main ones carried out approximately four months after the end of each program, and a shorter survey right after the end of the mentoring program from which we obtained information on salaries in jobs found during or right after the mentoring program. As discussed in Section 4, results based on the outcomes survey suffer from a selection problem. Thus, we interpret the findings based on this data as suggestive evidence.
@@
Figure 12 presents the share of respondents to mentoring outcomes surveys declaring a salary within a salary range (ranges were pre-defined in the survey). Orange bars indicate the treatment group, and the green ones are the control group. The right panel is based on the survey carried out right after the end of the program, while the left panel is from the survey conducted four months after the end of the program.
@@
Note: Reported salary ranges by subjects who found new jobs after the start of the mentoring program. Orange represents the treated group, and green represents the control group. The length of a bar represent the share of respondents declaring a salary within the range. The right panel survey was conducted right after the end of the program; the left panel survey was conducted four months after the program.

### v5

@@ 6.3 Counterfactual Mentors’ Selection Policies
Now, we examine selection within the Mentoring program: how do mentors choose mentees, and could alternative selection strategies improve program impact? Understanding selection matters for two reasons. First, econometrically, mentor selection determines the population for which our treatment effects are identified. Our estimates of $\tau^{M}(\mathcal{G}^{S})$ apply to the mentor-selected population, which may differ from the broader applicant pool. Second, mentoring occurs outside experimental settings, and mentor preferences may not align with social efficiency: if mentors favor applicants who would succeed regardless of intervention, or if less experienced mentors lack the judgment to identify high-potential candidates, there is scope for improved targeting. Mentors do demonstrably identify promising applicants: as shown in Figure 2, applicants selected by mentors but randomly assigned to the control group achieved substantially higher employment rates than non-selected applicants. But whether they identify applicants with high treatment effects—rather than high baseline employability—is an empirical question. Table 8 characterizes these selection patterns by comparing selected versus non-selected applicants (Panel A) and by comparing applicants chosen by different types of mentors (Panels B and C).
@@
Panel A reveals that mentors select applicants with substantially stronger baseline employment prospects: 46% of selected applicants in the control group find technology jobs, compared to 29% of non-selected applicants. Selected applicants are more likely to be over 30 (49% vs. 15%), hold STEM degrees (51% vs. 18%), and come from urban areas outside small towns (25% vs. 48% from small towns). This selection pattern presents a puzzle: Section 5.3 documented larger treatment effects for applicants from small towns and those without STEM degrees. Mentors appear to select on baseline employability rather than on treatment effect magnitude.
@@
Panels B and C suggest that experienced mentors employ a different selection philosophy. Mentors with managerial experience select applicants with weaker baseline employment prospects (38% vs. 55%) but who are older (59% vs. 38% over 30) and have more professional experience (7.6 vs. 6.5 years). Panel C shows similar patterns for mentors with above-median career experience. These experienced mentors appear to identify applicants with high potential who face significant barriers; perhaps recognizing that such applicants offer greater scope for mentoring impact. Less experienced mentors select more conventionally, favoring applicants with stronger baseline employability.
Alternative selection strategies.
We use our policy evaluation framework to examine how alternative selection strategies would affect Mentoring effectiveness. We consider four algorithmic approaches, each selecting 15% of applicants (approximating Dare IT’s typical capacity):
  • Highest Treatment Effect: select applicants with the largest predicted treatment effects $\hat{\tau}_{i}^{M}$
  • Most Promising: select applicants with the highest predicted outcomes absent treatment, $\hat{Y}_{i}(0)$
  • Least Promising: select applicants with the lowest predicted outcomes absent treatment
  • Best if Treated: select applicants with the highest predicted outcomes under treatment, $\hat{Y}_{i}(M)$
We benchmark these against random selection and the current mentor-based allocation.
Overlap between selection strategies.
Table 9 examines the extent to which different selection strategies choose the same individuals. Each cell shows the percentage of individuals selected under the row policy who also appear in the column policy. Importantly, the algorithmic policies do not use information on whether an applicant was actually selected by mentors; any overlap reflects convergence on similar applicant characteristics through independent criteria.
Table 9: Overlap Between Selection Strategies (%)
Highest TE Most Prom. Least Prom. Best if Treated Current
Highest Treatment Effect 100.0
@@
Best if Treated 27.9 26.7 31.5 100.0
Current (Mentor) 70.4 29.6 18.3 19.7 100.0
Notes: Each cell shows the percentage of row-policy selections that also appear in the column policy. Lower triangle only; upper triangle symmetric. All policies select 15% of applicants. Under random selection, expected overlap would be 15%.
The current mentor-based selection exhibits substantial overlap with the Highest Treatment Effect policy: 70% of mentor-selected applicants would also be selected by the treatment-effect-maximizing algorithm. This overlap far exceeds the 15% expected under random selection, indicating that mentors, through whatever criteria they apply, tend to choose applicants whom the algorithm also identifies as high-benefit. The overlap with Most Promising is considerably lower (30%), and overlap with Least Promising is lower still (18%), suggesting mentors do not primarily select based on baseline employability in either direction.
These patterns are consistent with two interpretations. First, mentors may implicitly target treatment effects, recognizing which applicants would benefit most from mentoring. Second, the characteristics that predict high treatment effects (such as facing network barriers) may correlate with other attributes mentors find compelling in application materials. We cannot distinguish these explanations, but the high concordance between mentor judgment and algorithmic targeting suggests that scaling the Mentoring program with algorithmic selection would likely preserve much of the effectiveness achieved under mentor-based selection.
Comparing selection strategies.
Table 10 compares treatment effects across six selection strategies for the Mentoring program. The Highest Treatment Effect policy achieves the largest impact: 3.59-3.59 months (SE 0.03), meaning selected participants find technology jobs 3.6 months faster than they would without the program. Random selection serves as a benchmark, generating effects of 2.20-2.20 months (SE 0.04). The Current mentor-based selection performs well, achieving 3.25-3.25 months (SE 0.10), a 48% improvement over random allocation. This validates that mentors successfully identify high-benefit applicants through their selection process.
Table 10: Treatment Effects Under Alternative Selection Strategies: Mentoring
Strategy Baseline Treatment Effect
@@
Algorithmic strategies:
Highest Treatment Effect 11.08 (0.09) -3.59 (0.03)
@@
Best if Treated 11.77 (0.15) -2.14 (0.08)
Notes: Comparison of selection strategies for Mentoring, each selecting 15% of applicants. “Baseline” is predicted months to employment without treatment (lower = faster job finding absent intervention). “Treatment Effect” is the change in months from treatment (negative = faster job finding due to program). Standard errors in parentheses.
The Highest Treatment Effect policy achieves a 10% improvement over mentor selection (3.59-3.59 versus 3.25-3.25 months). Notably, this algorithmic policy uses only observable application characteristics, while mentors access richer information including video introductions and holistic assessment. The substantial overlap between mentor selections and algorithmic targeting (70%, from Table 9) suggests that mentors and the algorithm identify similar applicant profiles through different information sets. Data-driven selection could complement mentor judgment by providing a systematic check or by scaling selection when mentor capacity is limited.
Policies targeting objectives other than treatment effects perform substantially worse. Most Promising selects applicants with the strongest baseline prospects—those predicted to find jobs in 9.7 months even without intervention—yet achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation. Strong baseline prospects do not predict strong treatment response. Conversely, Least Promising targets applicants with the weakest baselines (13.4 months to employment without intervention) but generates the smallest treatment effects (1.94-1.94 months). This pattern reveals a tension between equity and efficiency: applicants with the greatest baseline disadvantage benefit less from the program than those with moderate baselines. However, even the Least Promising group experiences meaningful gains: nearly two months of accelerated job finding, suggesting that equity-focused selection remains beneficial in absolute terms, if not maximal in efficiency terms.
Best if Treated, which selects applicants predicted to have the best outcomes under treatment rather than the largest treatment effects, performs similarly to random selection (2.14-2.14 versus 2.20-2.20 months). This underscores a key distinction: applicants who perform best after treatment are not the same as applicants who benefit most from treatment. Program impact depends on the counterfactual, what would have happened absent intervention, not just realized outcomes.
@@ 7 Conclusion
Changing labor markets push workers to want to transition into high-growth sectors, and facilitating these transitions is an important goal for labor policy. However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity. In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector, identifying a class of interventions that can effectively narrow the gender gap in technology.
Both programs generate large, statistically significant effects on technology employment. Mentoring increases technology employment by 15 percentage points (from 46% to 61%), while Challenges increases it by 11 percentage points (from 27% to 38%), with effects persisting throughout our observation period. The Challenges program operates at approximately \$15 per participant, implying a cost of roughly \$150 per additional technology job—a remarkably low figure relative to traditional training programs.
The two programs address distinct barriers to labor market entry. Mentoring expands professional networks: participants gain 53 LinkedIn connections on average, a 23% increase, while Challenges participants show no network growth. Challenges instead provides credible skill signals through portfolio projects that demonstrate technical competence to prospective employers. These distinct mechanisms generate predictable heterogeneity in treatment effects. Expanding professional networks creates the greatest opportunity for women from small towns, those without STEM degrees, and mothers. Building credible skill signals offers the greatest opportunity for early-career women with limited work history. The 28 percentage point larger effect for mothers in Mentoring is particularly striking, suggesting that mentoring can substantially reduce motherhood penalties in technology employment.
These heterogeneous effects have implications for program design. We develop an assignment policy that allocates applicants to their highest-benefit program based on predicted treatment effects. At current capacity levels, targeted assignment achieves effects 86% larger than random allocation. Even at expanded capacity where every applicant receives a program slot, targeting generates 23% larger effects than random assignment by matching individuals to their best-suited program. This demonstrates that the value of targeting comes not only from selecting high-benefit individuals but also from exploiting complementarities between participant characteristics and program mechanisms.
We also examine mentor-based selection within the Mentoring program. Mentors successfully identify promising candidates: their selections achieve effects 48% larger than random allocation. Algorithmic selection based on observable characteristics achieves a further 10% improvement. The 70% overlap between mentor and algorithmic selections suggests both approaches identify similar applicant profiles. Data-driven tools could complement mentor judgment, particularly for scaling selection when mentor capacity is limited.
Several limitations merit acknowledgment. First, moderate sample sizes yield wide confidence intervals for some estimates, particularly for subgroup analyses. Second, while our evidence is consistent with network and signaling mechanisms, we cannot definitively isolate the causal channels through which each program operates. Understanding these mechanisms more precisely could inform efforts to scale Mentoring by identifying its essential components. Third, for Challenges, our estimates come from the first program cohort; subsequent cohorts may differ in composition or in treatment response as the programs mature and selection patterns evolve. Yet studying small, custom-designed programs enabled the access and stakeholder cooperation necessary for our research design, including the novel paired randomization within mentor-selected candidates that allows us to separately identify selection and treatment effects, a contribution absent from prior mentoring RCTs.
More broadly, having two randomized interventions targeting distinct mechanisms within the same population enables comparisons rarely possible in program evaluation. We observe not only employment outcomes but also professional network expansion, allowing us to directly test whether mentoring operates through networks. We examine mentor selection behavior—how mentors choose, whether experienced mentors select differently, and how mentor choices compare to algorithmic efficiency. These analyses were possible only because our design preserves mentor discretion while enabling randomization.
Our findings demonstrate that women’s underrepresentation in technology reflects multiple distinct barriers—network access, credentialing, confidence—that require different interventions. Effective policy should provide multiple pathways tailored to different populations rather than assuming a single intervention serves all. The substantial heterogeneity we document, and the gains from exploiting it through targeted assignment, suggest that personalized approaches to workforce development may substantially outperform one-size-fits-all programs.
@@ Appendix B Covariate balance
In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only variable with a standardized absolute mean difference above the 0.1 threshold is is social science.
@@
Appendix C Monthly Outcomes Plots
Monthly Employment Outcomes.
Figure 7 presents proportions of subjects across programs and treatment groups who reported a job in the technology sector with the inclusion of subjects who reported a job without a month in which they started. When a subject reported a job without a month, but only with the year in which they started we assign the last month in the data.
Figure 7: Share of Applicants with a Tech Job per Group Over Time.
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we present outcomes of three Mentoring groups: the treatment group (in blue), the control group (in green), and the applicants who were not selected by mentors (in red). In the right figure, we group focus on the Challenges program groups. The average outcome in the treated group (in blue), the control group (in green) and the late applicants to the program who were not included in the randomization and the main analysis (in red).
We notice that there is an increase in the shares that are employed in the last month and this is particularly pronounced for non-selected Mentoring applicants.
Kaplan Meier Survival Curves.
@@ Appendix E Outcomes measured in surveys
We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes. There were three surveys: one main survey for each program, both administered about four months after completion, and a shorter survey right after the end of the Mentoring from which we obtained information on salaries in jobs found during or right after the mentoring program. As discussed in Section 4, results based on the outcomes survey suffer from a selection problem. Thus, we interpret the findings based on this data as suggestive evidence.
@@
Figure 13 presents the share of respondents to mentoring outcomes surveys declaring a salary within a salary range (ranges were pre-defined in the survey). Orange bars indicate the treatment group, and the green ones are the control group. The left panel is based on the survey carried out right after the end of the program, while the right panel is from the survey conducted four months after the end of the program.
@@
Note: Reported salary ranges by subjects who found new jobs after the start of the mentoring program. Orange represents the treated group, and green represents the control group. The length of a bar represent the share of respondents declaring a salary within the range. The left panel survey was conducted right after the end of the program; the right panel survey was conducted four months after the program.